# Boosting Adversarial Transferability with Low-Cost Optimization via Maximin Expected Flatness

### v2

@@ VI-A Experimental Setup
Inc-v3adv Inc-v3ens3 Inc-v3ens4 IncRes-v2ens RS HGD NRP Inc-v3adv Inc-v3ens3 Inc-v3ens4 IncRes-v2ens RS HGD NRP
MI [21] 21.6 17.2 16.4 9.4 25.6 19.4 53.5 27.3 22.4 23.3 11.2 24.5 8.9 30.7
NI [22] 21.8 18.3 17.3 9.2 25.7 17.0 55.4 27.8 22.8 22.6 11.1 24.4 8.4 30.5
PI [4] 22.7 19.1 17.5 9.8 26.6 21.6 60.7 30.9 24.5 24.3 12.6 24.7 9.8 30.9
TPA [2] 25.7 18.1 15.5 9.3 28.4 6.7 58.2 27.8 18.0 16.2 8.6 25.8 2.3 33.2
GNP [5] 29.2 24.8 23.2 13.3 28.1 26.5 70.7 30.5 24.0 23.5 11.7 25.1 6.5 34.4
VMI [1] 36.5 35.4 34.3 21.8 30.6 45.8 73.7 45.3 41.2 41.6 25.1 27.3 24.5 38.5
VNI [1] 37.8 35.5 34.7 23.8 30.9 47.4 75.9 47.8 44.4 43.9 26.8 28.0 26.7 38.9
EMI [4] 29.5 23.6 22.4 12.5 29.0 28.9 71.4 39.8 33.5 31.4 17.7 27.2 13.5 38.3
RAP [7] 29.4 17.4 16.9 8.8 32.1 4.0 63.2 32.2 16.3 16.2 7.9 27.3 1.5 27.6
APP [6] 51.1 50.0 49.5 35.6 39.6 61.3 89.0 53.7 47.2 47.3 29.9 31.4 24.1 44.2
FEM [3] 50.2 49.3 46.4 31.4 36.4 72.2 89.5 56.2 49.7 49.4 28.5 29.9 23.5 43.0
PGN [8] 60.2 58.4 59.6 45.6 48.1 62.5 92.9 69.2 61.1 62.2 41.6 38.1 27.7 54.6
MEFH 62.4 64.5 63.3 50.4 49.2 80.6 94.6 69.7 65.1 65.4 45.4 38.3 46.3 54.8
MEFF 70.1 69.2 65.1 52.2 51.5 83.0 97.2 75.8 69.0 69.6 47.9 40.4 52.4 56.2
TABLE IV: Transfer attack success rate (%) comparison with input augmentation attacks. Using Res-50 and Inc-v3 as surrogates, MEFH/F variants outperform five input augmentation baselines across six targets while using less computation. The best results are bold and the second best results are underlined.
Attack Res-50 $\Longrightarrow$ Inc-v3 $\Longrightarrow$
IncRes-v2 Inc-v3ens3 Inc-v4ens4 IncRes-v2ens ViT-L/32 MLP-Mixer IncRes-v2 Inc-v3ens3 Inc-v4ens4 IncRes-v2ens ViT-L/32 MLP-Mixer
DI [19] 70.2 37.5 36.5 23.1 29.5 59.2 66.7 37.1 35.7 18.8 26.3 54.3
MEFH+DI 89.7 77.5 76.1 66.1 50.3 74.5 88.3 74.8 74.7 56.2 41.6 67.8
MEFF+DI 95.3 83.8 80.5 68.6 55.1 79.8 94.2 78.0 77.6 58.0 46.4 74.1
TI [20] 51.1 42.0 39.1 30.2 34.4 55.6 52.0 37.0 37.2 26.0 28.8 53.7
MEFH+TI 81.7 77.4 77.8 68.2 57.3 71.1 86.4 79.6 79.5 67.4 48.7 68.3
MEFF+TI 94.1 83.5 82.6 73.6 62.5 76.0 94.5 89.0 88.3 78.8 56.3 75.6
SI [22] 56.8 29.6 29.5 17.4 24.9 49.5 65.2 34.7 35.3 18.3 25.1 53.7
MEFH+SI 86.9 71.3 69.7 56.8 44.1 67.1 89.4 75.7 75.0 55.2 39.4 64.2
MEFF+SI 92.8 78.6 74.7 61.2 48.8 73.0 95.9 81.7 80.5 57.8 44.9 72.2
Admix [25] 52.5 23.6 22.2 13.1 23.9 50.4 64.2 29.7 27.7 14.7 23.6 52.8
MEFH+Admix 70.5 39.2 37.0 24.5 27.0 55.5 83.5 45.3 45.8 27.1 26.5 56.5
MEFF+Admix 86.4 46.6 42.1 27.7 31.6 64.1 94.7 51.1 51.9 30.1 31.0 67.6
SSA [26] 83.0 50.2 47.3 31.3 35.6 64.7 83.7 52.4 54.7 31.0 33.3 62.3
MEFH+SSA 81.6 53.9 51.1 38.0 33.0 61.1 84.4 54.9 56.3 34.2 30.9 58.2
MEFF+SSA 90.1 58.1 53.5 39.0 35.7 68.0 92.2 58.2 60.3 35.1 35.3 65.7
MEFH 83.7 64.5 63.3 50.4 40.8 65.3 86.1 65.1 65.4 45.4 35.2 65.5
MEFF 90.3 69.2 65.1 52.2 44.9 71.7 94.3 69.0 69.6 47.9 40.3 69.4
VI-B Evaluation on Normally Trained Models
We evaluate our MEFH and MEFF attacks against 12 representative gradient stabilization methods on normally trained models. For each baseline, we generate adversarial examples on four standard source models (Res-50 [9], Res-101 [9], Inc-v3 [48], and Inc-v4 [57]), then measure attack success rates across eight diverse target models: Inc-v3 [48], Inc-v4 [57], Res-50 [9], Res-101 [9], IncRes-v2 [57], Dense-121 [59], VGG-19 [60], and Xception [61].
As shown in Table II, both MEF variants significantly outperform existing methods. Notably, MEFH achieves an average attack success rate of 89.93% across all targets while requiring only half the computational budget of PGN [8]. This surpasses PGN’s 88.18% average success rate despite using 50% fewer backpropagation steps. When matching PGN’s computational cost, MEFF elevates the average success rate to 94.84%, a 6.66% absolute improvement over PGN. Remarkably, MEFF maintains over 90% success rates on every individual target model, demonstrating unprecedented cross-model transferability. These results establish that our MEF attack effectively balances efficiency and attack performance, achieving new SOTA transferability under both half-cost and full-cost settings.
VI-C Evaluation on Defended Models
To evaluate robustness against modern defenses, we generate adversarial examples on Inc-v3 [48] and Res-50 [9] source models, then transfer them to seven protected targets: (1) adversarially trained models (Inc-v3adv [37], Inc-v3ens3 [38], Inc-v3ens4 [38], IncRes-v2ens [38]), (2) purification defenses (HGD [41], NRP [43]), and (3) certifiable adversarial defense RS [39]. For NRP [43], we evaluate attack success rates on Res-101 [9] after applying its purification process.
@@ VI-C Evaluation on Defended Models
Table III compares MEFH/MEFF with 12 gradient stabilization attacks. MEFH achieves 59.87% average success rate across all defenses using half the computational budget of PGN [8], surpassing PGN’s 55.84% baseline. When matching PGN’s cost, MEFF further elevates the average success rate to 62.89%, a 7.05% absolute improvement. Notably, MEFF outperforms all competitors on individual defense types: 7.63% against adversarial training, 12.73% over purification defenses, and 4.1% versus RS certification. This consistent superiority demonstrates our method’s capability to bypass diverse defense mechanisms, establishing new state-of-the-art robustness against protected models.

### v3

@@ VI-B2 Evaluation on Diverse Network Architectures
[…]
VI-B3 Attacking on Ensemble of Models
We extend our evaluation to ensemble attacks by integrating Inc-v3 [37] and Res-50 [9] through logit averaging as the surrogate model. Table IV compares MEF against the top-4 performing baselines (full comparison with 14 methods, including ensemble-specialized SVRE [24] and CWA [25], is provided in Supp. -H). MEFF demonstrates consistent superiority, outperforming the state-of-the-art PGN [8] by margins of 1.6%-8.2% across diverse targets. This confirms that MEF captures transferable adversarial patterns across model ensembles.
VI-B4 Integrated with Input Augmentation Attacks
We compare MEFH and MEFF against 10 representative input augmentation attacks (e.g., SSA [19], DeCoWA [74]). Following the standard setup, we generate adversarial examples on Res-50 [9] and Inc-v3 [37] and evaluate their transferability to challenging targets, including ensemble models and heterogeneous architectures. Furthermore, to demonstrate compatibility, we integrate sample transformations from five classical augmentation methods (e.g., TI [15], DI [14]) into MEF variants without introducing additional gradient computations.
[…]
TABLE VI: Transfer attack success rate (%) comparison across attack paradigms. The best results are bold.
SGM [22] 71.6 35.3 92.8 12.3 12.5 6.1 38.4
LinBP [23] 72.7 47.9 95.4 13.2 13.3 6.8 41.6
AWT [76] 89.1 87.5 97.9 63.1 61.4 49.1 74.7
FAUG [77] 68.0 64.1 78.6 41.5 42.2 27.4 53.6
ANA [78] 79.8 75.4 86.0 51.4 50.3 38.9 63.6
LL2S [79] 88.2 86.6 96.7 61.9 60.6 47.8 73.6
ILA [26] 70.5 74.8 97.9 16.5 15.3 9.0 47.3
NAA [27] 82.7 82.4 98.3 30.1 29.6 21.0 57.4
BFA [80] 89.1 87.3 98.6 63.2 60.9 48.6 74.6
P2FA(10,70) [81] 95.4 95.7 99.8 70.1 66.4 54.2 81.5
CDA [28] 69.4 83.1 94.5 49.8 52.7 43.3 65.5
DiffAttack [82] 41.4 41.9 52.6 15.9 14.0 5.4 28.5
MEFH 90.6 88.7 99.7 64.5 63.3 50.4 76.2
MEFF 95.1 94.2 99.9 69.2 65.1 52.2 79.3
MEF(10,70) 97.5 96.6 99.9 73.1 68.5 55.7 83.6
VI-B5 Comparison with Other Types of Transfer-based Attacks
We validate MEF against three paradigms: surrogate refinement (SGM [22], LinBP [23], AWT [76], FAUG [77], ANA [78], LL2S [79]), feature-based (ILA [26], NAA [27], BFA [80], P2FA [81]), and generative attacks (CDA [28], DiffAttack [82]). As shown in Table VI, MEF consistently outperforms these approaches on Res-50 [9]. Notably, MEFF (79.3%) surpasses leading baselines like AWT [76] (74.7%) and BFA [80] (74.6%) solely through geometry-aware input space optimization, avoiding reliance on architectural priors.
Among feature-based methods, the state-of-the-art method P2FA [81] attains 81.5% but incurs $3.5\times$ higher computational cost. Under a budget-aligned setting, our MEF(10,70) variant reclaims the lead with 83.6% (+2.1% over P2FA), confirming the fundamental superiority of zeroth-order flatness optimization over intermediate feature manipulation.
TABLE VII: Comparison of Efficiency, Constraint Compliance, and Stealthiness on Inc-v3 ($\epsilon=16/255$).
Method Category Queries ASR (%) LPIPS $\downarrow$ SSIM $\uparrow$
Boundary [12] Decision-based 500 12.3 0.85 0.291
Boundary [12] Decision-based 5,000 24.5 0.72 0.346
RayS [84] Decision-based 500 59.8 0.18 0.887
RayS [84] Decision-based 5,000 94.3 0.05 0.947
SimBA [13] Score-based 500 7.80 0.01 0.991
SimBA [13] Score-based 5,000 41.8 0.01 0.992
Square [83] Score-based 500 94.9 0.31 0.583
Square [83] Score-based 5,000 99.7 0.30 0.579
MEFH Transfer-based 0 90.6 0.24 0.751
MEFF Transfer-based 0 95.1 0.30 0.692
VI-B6 Comparison with Query-based Attacks
We benchmark MEF against score-based (SimBA [13], Square [83]) and decision-based (Boundary [12], RayS [84]) attacks on Inc-v3 [37]. Unlike feedback-dependent baselines, MEF operates in a zero-query setting using a ResNet-50 surrogate. As shown in Table VII, while Square Attack [83] achieves marginally higher success (99.7%), it incurs prohibitive costs (5,000 queries) and severe visual degradation (SSIM 0.579). In contrast, our zero-query $\text{MEF}_{\text{F}}$ attains a competitive 95.1%, outperforming the decision-based SOTA RayS [84] (94.3% at 5,000 queries). Even the half-cost $\text{MEF}_{\text{H}}$ maintains 90.6% success with superior stealth (SSIM 0.751). This confirms MEF offers a practical solution without the detection risks and computational overhead of high-frequency queries.
VI-C Practical Threats to Real-World Applications
TABLE VIII: Transfer Attack Success Rates on Six Commercial Vision APIs. The best results are bold and the second best are underlined.
Method Google Cloud Vision Amazon Rekognition Azure AI Vision
T-1 T-5 T-10 T-1 T-5 T-10 T-1 T-5 T-10
PGN [8] 67.20% 44.30% 40.00% 54.50% 36.60% 26.30% 59.10% 32.00% 27.40%
FEM [3] 65.30% 44.10% 39.00% 57.10% 40.00% 31.00% 54.40% 26.10% 21.20%
FGSRA [45] 68.60% 46.10% 40.90% 55.90% 39.70% 29.20% 51.70% 22.30% 18.50%
MEFH 71.60% 50.20% 46.10% 58.50% 42.80% 33.90% 60.90% 34.70% 29.10%
MEFF 79.40% 63.30% 59.00% 68.80% 55.50% 46.70% 67.18% 40.33% 34.88%
Method Baidu AI Cloud Alibaba Cloud Tencent Cloud
T-1 T-5 T-10 T-1 T-5 T-10 T-1 T-5 T-10
PGN [8] 79.80% 61.10% 61.10% 86.40% 66.10% 66.10% 74.90% 53.10% 51.90%
FEM [3] 80.60% 61.30% 61.30% 83.40% 64.20% 64.20% 72.50% 50.80% 49.80%
FGSRA [45] 74.50% 57.80% 57.80% 83.00% 66.00% 66.00% 76.00% 52.70% 51.50%
MEFH 83.40% 62.70% 62.70% 86.80% 66.80% 66.80% 75.10% 54.70% 53.60%
MEFF 86.30% 70.10% 70.10% 91.70% 77.30% 77.30% 84.70% 67.10% 66.10%
VI-C1 Evaluation on Commercial Vision APIs
To assess practical threats beyond open-source models, we evaluate transferability against six commercial black-box systems. Unlike local models, these APIs run proprietary architectures with unknown preprocessing defenses. Given query costs, we compare MEF against the three leading baselines using the Top-k Exclusion (T-k) metric, where the correct ground-truth label is excluded from the top-k predictions.
Table VIII demonstrates that MEF consistently outperforms baselines across all platforms. Notably, on the rigorous Google Cloud Vision API, MEFF achieves 79.40% T-1, surpassing the SOTA PGN (67.20%) by a significant margin of 12.2%. Even the half-cost MEFH (71.60%) exceeds the strongest baseline. High success rates are also observed on Alibaba Cloud (91.70%) and others. These results confirm that the flat local minima identified by our framework generalize effectively to unknown production pipelines. For intuitive visual evidence of these successful semantic subversions, we provide qualitative comparisons of API outputs in Fig. A-4 of the Supp. -L.
TABLE IX: Transferability comparison (Fooling Rate %) against Multimodal Large Models. The best results are bold and the second best are underlined
Attack Method LLaVA Unidiffuser MiniGPT-4 Qwen2.5-VL Average
APP [6] 19.1 24.9 23.8 23.4 22.8
FEM [3] 22.5 29.3 26.5 29.4 26.9
PGN [8] 22.0 32.4 27.8 24.5 26.7
FGSRA [45] 20.3 26.4 25.7 24.7 24.3
ANDA [44] 19.6 25.5 24.9 23.9 23.5
GI [46] 19.8 25.7 25.0 24.1 23.6
MUMODIG [47] 20.2 26.2 25.5 24.5 24.1
GAA [48] 19.8 25.8 25.1 24.1 23.7
FoolMix [49] 19.4 25.2 24.5 23.6 23.2
MEFH 26.7 35.4 30.5 33.4 31.5
MEFH 30.5 38.4 35.5 41.5 36.5
VI-C2 Transferability to Multimodal Large Language Models
To investigate transferability against Multimodal Large Language Models (MLLMs), we evaluate MEF on LLaVA [68], Unidiffuser [69], MiniGPT-4 [70], and Qwen2.5-VL [71]. We generate adversarial examples on Res-50 [9] and transfer them to these models. This challenging cross-task scenario requires disrupting semantic alignment across modalities.
We prompt models with “Describe the image” to generate captions $C_{cln}$ and $C_{adv}$ for clean and adversarial images. Let $D(\cdot)$ denote the CLIP text encoder [85]. We report the CLIP-based Fooling Rate, which is defined as the proportion of samples satisfying the semantic divergence condition:
$\text{CosSim}\big(D(C_{cln}),D(C_{adv})\big)\leq\tau,$ (15)
where $\tau=0.4$ is an empirical threshold adopted to distinguish successful attacks from benign caption variations.
As shown in Table IX, MEF demonstrates superior cross-paradigm transferability. MEFF achieves the highest average fooling rate of 36.5% across all models. Notably, even the efficient MEFH (31.5% average) outperforms the best baseline FEM (26.9%). On Qwen2.5-VL, MEFF reaches 41.5%, surpassing baselines by over 12%. This confirms that optimizing multi-order flatness creates robust adversarial features capable of compromising the visual encoders of advanced MLLMs (see Supp. -I for qualitative examples).
VI-D Robustness against Defense Mechanisms
TABLE X: Transfer attack success rate ($\%$) against modern defense mechanisms. Evaluating across three defense categories on seven protected targets, our MEFH/F outperform all competitors. The best results are bold and the second best are underlined.
@@ VI-D Robustness against Defense Mechanisms
Inc-v3adv Inc-v3ens3 Inc-v3ens4 IncRes-v2ens RS HGD NRP AVG Inc-v3adv Inc-v3ens3 Inc-v3ens4 IncRes-v2ens RS HGD NRP AVG
APP [6] 50.1 49.0 48.5 34.6 38.6 60.3 88.0 52.7 52.7 46.2 46.3 28.9 30.4 23.1 43.2 38.7
FEM [3] 50.2 49.3 46.4 31.4 36.4 72.2 89.5 53.6 56.2 49.7 49.4 28.5 29.9 23.5 43.0 40.0
PGN [8] 60.2 58.4 59.6 45.6 48.1 62.5 92.9 61.0 69.2 61.1 62.2 41.6 38.1 27.7 54.6 50.6
ANDA [44] 50.2 48.9 47.3 37.7 41.0 62.5 87.7 53.6 58.2 50.2 50.7 29.5 29.4 22.3 45.1 40.8
FGSRA [45] 52.8 51.4 52.1 36.7 39.8 63.6 91.6 55.4 57.9 50.5 51.1 31.9 30.5 26.1 45.2 41.9
GI [46] 50.5 50.9 48.1 37.1 37.9 60.7 89.7 53.6 57.5 51.3 52.2 32.0 32.1 22.5 44.1 41.7
MUMODIG [47] 53.2 51.6 51.8 36.9 40.2 63.7 87.7 55.0 57.2 51.0 52.1 32.2 32.1 23.0 46.1 42.0
GAA [48] 49.7 49.0 51.1 36.9 39.5 62.8 85.6 53.5 55.2 49.7 50.0 31.5 31.7 25.0 45.1 41.2
FoolMix [49] 49.6 50.7 46.2 34.9 38.1 62.2 84.5 52.3 55.6 48.5 49.5 29.7 32.3 23.4 42.3 40.2
MEFH 62.4 64.5 63.3 50.4 49.2 80.6 94.6 66.4 69.7 65.1 65.4 45.4 38.3 46.3 54.8 55.0
MEFF 70.1 69.2 65.1 52.2 51.5 83.0 97.2 69.8 75.8 69.0 69.6 47.9 40.4 52.4 56.2 58.8
VI-D1 Evaluation on Standard Defense Methods
@@
To evaluate robustness, we transfer adversarial examples generated on Inc-v3 and Res-50 to seven representative protected targets spanning adversarial training [29, 30], purification [32, 33], and certifiable defense [31], as detailed in Sec. IV-A. Table X compares MEFH/MEFF with competitive gradient stabilization attacks that achieve an average success rate exceeding 40%. Results for other baselines (e.g., MI [16], VMI [1]) are provided in Supp. -F (Table A-3). MEFH achieves 60.70% average success rate across all defenses using half the computational budget of PGN [8], a 4.86% improvement over PGN’s 55.84% baseline. When matching PGN’s cost, MEFF further elevates the average success rate to 64.30%, a 8.46% absolute improvement. Notably, MEFF outperforms all competitors on individual defense types. This consistent superiority demonstrates our method’s capability to bypass diverse defense mechanisms, establishing new state-of-the-art robustness against protected models.
TABLE XI: Transfer attack success rate ($\%$) against advanced defenses. The best results are bold and the second best are underlined.
Attack Diffusion-based Purifier RobustBench Models
DiffPure[34] DensePure[35] MIMIR[65] ARES[66] MeanSparse[67]
 Surrogate Model: ResNet-50
APP [6] 40.2 40.0 8.1 9.1 9.9
FEM [3] 42.8 42.0 8.1 9.1 9.9
PGN [8] 46.9 46.7 9.5 10.7 11.6
ANDA [44] 39.3 39.1 8.0 8.9 9.7
FGSRA [45] 41.3 43.7 9.1 10.2 11.1
GI [46] 40.9 40.7 8.3 9.3 10.1
MUMODIG [47] 41.4 41.3 8.4 9.4 10.2
GAA [48] 39.4 39.2 8.0 9.0 9.7
FoolMix [49] 40.7 40.5 8.3 9.3 10.1
MEFH 44.5 44.3 8.8 9.8 10.7
MEFF 48.6 47.2 17.3 19.4 21.1
 Surrogate Model: Inception-v3
APP [6] 34.8 34.4 7.0 7.9 8.6
FEM [3] 41.6 40.9 8.4 9.4 10.2
PGN [8] 45.1 44.5 9.1 10.2 11.1
ANDA [44] 37.1 36.6 7.5 8.4 9.1
FGSRA [45] 44.9 44.2 8.7 9.8 10.6
GI [46] 37.9 37.4 7.6 8.6 9.3
MUMODIG [47] 37.6 37.1 7.6 8.5 9.3
GAA [48] 36.7 36.2 7.4 8.3 9.0
FoolMix [49] 35.8 35.3 7.2 8.1 8.8
MEFH 43.7 42.9 8.5 9.6 10.4
MEFF 46.1 45.1 15.8 17.8 19.3
VI-D2 Evaluation on Advanced State-of-the-Art Defenses
To address the rapid evolution of defenses, we extend evaluation to state-of-the-art mechanisms: diffusion-based purification (DiffPure [34], DensePure [35]) and top-performing robust models from RobustBench [64] (MIMIR [65], ARES [66], MeanSparse [67]). For comparison, we select the top-9 performing baselines from Table X. As shown in Table XI, these advanced defenses pose significant challenges. Nevertheless, MEFF consistently demonstrates superior resilience. Against purification defenses, it outperforms the runner-up PGN [8] by 1.7% on DiffPure [34]. More notably, against the highly robust MeanSparse model [67], MEFF achieves a remarkable success rate of 21.1% (ResNet-50 [9] as surrogate model), nearly doubling PGN [8]’s 11.6%. This suggests that MEF’s average-case flatness optimization generates adversarial examples that are structurally stable enough to survive stochastic purification and robust feature alignment.
VI-E Analysis of Efficiency, Stealthiness, and Sensitivity
VI-E1 Convergence Analysis and Iteration Selection
Figure 5: Attack Success Rate (ASR) convergence curve over 100 iterations (Source: Res-50 [9] $\rightarrow$ Target: Inc-v3 [37]). The dashed red line marks the conventional $T=10$ iteration cutoff.
TABLE XII: Convergence and Performance Comparison. $\Delta$ denotes the improvement from $T_{10}$ to $T_{max}$. The best results are bold.
Method $\mathbf{T_{10}}$ ASR (%) $\mathbf{T_{\text{max}}}$ ASR (%) $\mathbf{\Delta}$ ASR (%) $\mathbf{T_{\text{max}}}$ Iter.
MI [16] 50.50% 51.20% 0.70% 56
NI [17] 55.70% 56.60% 0.90% 12
PI [4] 61.00% 61.00% 0.00% 10
FEM [3] 89.50% 89.60% 0.10% 12
FGSRA [45] 87.10% 89.70% 2.60% 23
PGN [8] 86.00% 88.50% 2.50% 40
MEF (Ours) 89.80% 96.40% 6.60% 36
To evaluate the impact of iteration count $T$ on transferability, we conducted an extended convergence analysis ($T=100$) using Res-50 [9] as the surrogate and Inc-v3 [37] as the target. We selected six representative baselines, comprising both classic and SOTA methods. Since attackers in realistic black-box scenarios lack target model feedback to identify the optimal stopping point, assessing whether attacks can sustain transferability over extended iterations is crucial.
The results, visualized in Fig. 5 and detailed in Table XII, reveal distinct behaviors. For most baselines, including the SOTA PGN [8], performance rapidly plateaus or even declines after reaching an early peak. This trend aligns with established findings [16], indicating that excessive iterations lead to overfitting on the surrogate model, thereby hampering transferability. Consequently, the gap between their $T_{10}$ performance and absolute peak ($T_{max}$) is marginal ($3\%$), validating $T=10$ as a cost-effective standard for these methods. In contrast, MEF effectively mitigates overfitting through flatness optimization, exhibiting a significantly higher optimization ceiling. It continues to accumulate gains up to $T=36$, achieving a peak ASR of 96.40% (a 6.6% improvement over $T=10$). Crucially, MEF remains strictly superior to all competitors throughout the entire 100-iteration process, confirming its dominance for both resource-constrained ($T=10$) and determined attackers.
VI-E2 Perceptual Quality Analysis
Figure 6: Comparison of Transferability (ASR) vs. Visual Quality (SSIM). Transfer ASR are valuated on Res-50 [9] $\to$ Inc-v3 [37] with $\epsilon=16/255$.
TABLE XIII: Quantitative comparison of visual perceptibility. The best results are bold and the second best are underlined.
Method Transferability Visual Quality Metrics
ASR (%) $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$
MUMODIG [47] 84.4 0.696 26.86 0.323 0.247
FGSRA [45] 85.3 0.683 26.62 0.313 0.252
PGN [8] 86.3 0.681 26.36 0.294 0.250
FEM [3] 89.4 0.666 26.44 0.322 0.251
MEFH 90.6 0.721 28.35 0.250 0.241
MEFF 95.1 0.692 26.56 0.295 0.244
We evaluate perceptual quality ($\epsilon=16/255$) via the ASR-SSIM trade-off (Fig. 6) and quantitative metrics (Table XIII). As shown in Fig. 6, MEF occupies the optimal top-right quadrant, balancing strength and stealth. Quantitatively, $\text{MEF}_{\text{H}}$ achieves the best fidelity (SSIM 0.721, LPIPS 0.250), surpassing the prior leader MUMODIG [47]. Meanwhile, $\text{MEF}_{\text{F}}$ yields the highest transferability (95.1%) with competitive stealth (outperforming PGN [8] in DISTS). This confirms MEF’s superiority stems from geometric efficiency. Qualitative results are in Supp. -J.
VI-E3 Computational Efficiency Analysis
TABLE XIV: Computational efficiency and resource cost comparison. We report computational cost (Backpropagations/BP), Peak GPU Memory (MiB), and average runtime (seconds/image).
Method Time (s/img) $\downarrow$ BP (count) $\downarrow$ Peak Mem. (MiB) $\downarrow$ ASR (%) $\uparrow$
MI [16] 0.29 10 1397 50.5
NI [17] 0.29 10 1400 56.5
PI [4] 0.30 10 1403 60.2
VMI [1] 0.56 200 4316 72.8
VNI [1] 0.59 200 4315 77.2
EMI [4] 0.48 200 4309 75.0
FEM [3] 2.80 210 1,788 89.4
FGSRA [45] 0.96 400 4100 85.3
PGN [8] 1.13 400 4488 86.3
MEFH 0.46 200 4090 90.6
MEFF 0.97 400 4097 95.1
We evaluate computational efficiency by measuring the hardware-agnostic Backpropagation (BP) counts, Peak GPU Memory, and the average wall-clock runtime required to generate an adversarial example (averaged over 100 randomly selected samples on an NVIDIA RTX 4090). For fair comparison, all sampling-based methods employ $N=20$ points with full synchronous parallel computation. We benchmark against basic single-point attacks and representative, well-established sampling-based methods.
As detailed in Table XIV, single-point methods are naturally the fastest (10 BPs, $\approx$ 1400 MiB) but suffer from limited transferability. In contrast, MEFH achieves a remarkable 90.6% success rate with only 200 BPs and 4090 MiB memory. Crucially, compared to the flatness-based SOTA PGN [8] (400 BPs, 4488 MiB), MEFH halves the gradient cost and reduces memory footprint, delivering a $\mathbf{2.4\times}$ speedup while improving ASR by 4.3%. Even our full-cost variant MEFF remains faster than PGN with lower memory usage, and significantly outpaces heavy methods like FEM [3] (2.80s). This confirms that our GBO strategy effectively mitigates the computational burden, achieving an optimal efficiency-performance ratio.